OpenStudy (welshfella):

My grandson asked me this one - but i need help! Integrate sin5x cos2x using a method other than integration by parts. Any ideas?

OpenStudy (amistre64):

use the trig identitys of cos addition and cosine subtraction

OpenStudy (amistre64):

what is cos(a+b) and cos(a-b)?

OpenStudy (amistre64):

bah .. sin cos is in sin addition and subtraction lol

OpenStudy (welshfella):

oh ok yeas i know what they are

OpenStudy (amistre64):

what is ... sin(a+b) and sin(a-b) those should work out better

OpenStudy (welshfella):

so use sin (a +b) and sin (a -b) ?

OpenStudy (amistre64):

type them for me and lets see how they are useful

OpenStudy (amistre64):

by the way, this is the impetus behind fourier series ...

OpenStudy (welshfella):

sin (a + b) = sina cosb + cosa sinb sin(a -b) = sina cosb - cosa sinb

OpenStudy (welshfella):

/oh ! add them together?

OpenStudy (amistre64):

yes ;) now when we add the 2 togheter, the rightest terms go to zero sin(a+b) + sin(a-b) = 2 sina cosb

OpenStudy (welshfella):

right i got it now plug in 5x and 2x for a and b

OpenStudy (amistre64):

very good

OpenStudy (welshfella):

and we'll get the sum of sin and cos

OpenStudy (welshfella):

clever stuff

OpenStudy (amistre64):

\[sin(a)cos(b)=\frac12[sin(a+b)+sin(a-b)]\] \[sin(5x)cos(2x)=\frac12[sin(7x)+sin(3x)]\]

OpenStudy (welshfella):

yea

OpenStudy (welshfella):

now its easy to integrate thanx

OpenStudy (amistre64):

youre welcome :)

OpenStudy (welshfella):

I'm not sure if i did this in the past. Must have i guess

OpenStudy (amistre64):

it wasnt covered that well in my courses, i only really learned it when i was going over fourier series. they use this techique to find the coefficients of their series.

OpenStudy (welshfella):

I did fourier series but its now so hazy lol!

OpenStudy (amistre64):

heheh, oh i know that feeling all to well

OpenStudy (welshfella):

yea ty